If tana = n tan Band sin A = m sin B, prove that Cos2 A = m2 – 1/ n2 – 1

Grade:10

1 Answers

Vijay Mukati
8 years ago
Hint: On simplification we get, cosA = mcosB/n.
Now since cos2A = 1-sin2A = 1-m2(1-cos2B) = 1 – m2 + n2cos2A.
On simplification, we will get cos2A = (m2-1)/(n2-1). Thanks